OpenStudy (anonymous):

I'm looking to determine the convergence or divergence of the geometric series E^(infinity)_ j =1. 4(-1/3)^j-1. *that is the autosome of j to infinity*

OpenStudy (anonymous):

I don't speak French, sorry.

OpenStudy (kc_kennylau):

Where are you from? I judged that by your name...

OpenStudy (anonymous):

The US. (Imported)

OpenStudy (kc_kennylau):

From?

OpenStudy (anonymous):

Mexico

OpenStudy (kc_kennylau):

Porque yo no entiendo lo que escribite tu...

OpenStudy (kc_kennylau):

Es \(\large\sum\limits_{j=1}^\infty \left[4(-\frac13)^j-1\right]\) lo que significas?

OpenStudy (anonymous):

That's the one! How do you enter graphics?

OpenStudy (anonymous):

I don't have the autosum key

OpenStudy (kc_kennylau):

So you don't speak Spanish? I learnt \(\LaTeX\) by myself, here is what I typed: \(\mbox{\ }\mbox{(}\) \sum_{j=1}^\infty \left[ 4(-\frac13)^j -1 \right] \(\mbox{\ )}\mbox{}\)

OpenStudy (anonymous):

That's awesome, thanks. As far as speaking Spanish, I'm second- generation import, and didn't learn a lot of spanish. I'm actually learning, now, but I understand it mostly.

OpenStudy (kc_kennylau):

I see. Try to learn \(\LaTeX\), it can help you a lot :) \ (\LaTeX\ ) --> \(\LaTeX\)

OpenStudy (anonymous):

\ (\sum_{j=1}^1\infty\left[4(-frac13)^j-1\right]\)

OpenStudy (anonymous):

(needs practice )

OpenStudy (kc_kennylau):

so when i'm displaying the code I would leave a gap between \ and ( but when you really need to type the code out, don't.

OpenStudy (kc_kennylau):

By the way, \ and ( is for in-line, while \ and [ will give you a newline

OpenStudy (anonymous):

Thanks for these tips!

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (kc_kennylau):

So you're in an English-speaking country, with a Spanish-speaking origin, and a French surname?!

OpenStudy (kc_kennylau):

Wait, is Ramosville your surname

OpenStudy (anonymous):

Haha, all of the Ramos-usernames were taken. So I went with Ramos-ville. Its from the song Margaritaville

OpenStudy (kc_kennylau):

oh lol

OpenStudy (anonymous):

Sweet, thanks for the medal!

OpenStudy (kc_kennylau):

\[\lim_{j\rightarrow\infty}\frac{\left[ 4(-\frac13)^{j+1} -1 \right]}{\left[ 4(-\frac13)^j -1 \right]}\] \ [\lim_{j\rightarrow\infty}\frac{\left[ 4(-\frac13)^{j+1} -1 \right]}{\left[ 4(-\frac13)^j -1 \right]}\]

OpenStudy (anonymous):

So this formula is what is used to determine convergence?

OpenStudy (anonymous):

What is your background in math? You seem pretty sharp

OpenStudy (kc_kennylau):

yep, and my father taught me much about math

OpenStudy (anonymous):

So you teach, or do this for fun?

OpenStudy (kc_kennylau):

Look at my profile, I'm just 14 years old, how can i teach lol

OpenStudy (anonymous):

Wow, that is amazing. You are making your father proud.

OpenStudy (kc_kennylau):

Gracias :)

OpenStudy (kc_kennylau):

Y tu, cuantos anos tienes?

OpenStudy (anonymous):

I am old! Hey, I had one quick question ! How did you arrive at that formula?

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Ratio_test#The_test

OpenStudy (anonymous):

Great, thank you for all your help. You have a bright future, young friend.

OpenStudy (kc_kennylau):

No problem and thank you :)